OpenStudy (anonymous):

Alex the plumber requires 6 lengths of cooper tubing to finish a fire sprinkler system. The length required are 3 at 2 feet, and 3 at 16 inches. How much will be leftover for the next job if he starts with a 16 foot long tube?

OpenStudy (cwrw238):

first work out the total length of the tubes he uses = 3 * 2 + 3 * 1 ft 4 ins

OpenStudy (cwrw238):

then subtract this from 16

OpenStudy (ribhu):

first of all 16 inches = 4/3 feets (since 12 inces = 1 feet) so the total length would be 3*2 (feet) = 6 feets + 3*4/3 = 4 feets total becomes 10 feets so the pipe remaining would be 6 feets

OpenStudy (cwrw238):

hmm...

OpenStudy (anonymous):

nice!

OpenStudy (ribhu):

@katlin95 did u understand it thoroughly?

OpenStudy (anonymous):

yes i did!

OpenStudy (cwrw238):

@ribhu - I'm glad you asked that last question